Fermi National Laboratory


Accelerator Update for January 11, 2006 - January 13, 2006

Wednesday January 11

The day shift began with Operations monitoring store #4587, the Recycler Ring (RR) stash at 193.4 E10, Pbar stack at 65.8 E10. Beam to NuMI, MiniBooNE is off for access, MCenter is off for analyzing magnet repair.

Operations put the Recycler into shot setup at 9:02 AM to stash antiprotons.

Pbar returns to stacking at 9:52 AM.

Tevatron (Tev) end of store studies began at 2:06 PM.

Operations put the Recycler into shot setup at 3:45 PM to stash antiprotons.

At 4:13 PM, the temperature problem with MCenter quad MC6Q5 has been resolved.

Pbar returns to stacking at 4:42 PM.

Store #4587 was terminated at 6:05 PM.

MCenter turned back on at 7:12 PM, they are running without their analyzing magnets. Work will continue on MC7AN1 tomorrow.

Shot setup begins at 7:55 PM.

Operations established store 4589 at 11:04 PM with an initial luminosity of 114.772 E30.

Thursday January 12

The midnight shift began with Operations monitoring store #4589, the Recycler Ring (RR) stash at 7.6 E10, Pbar stack at 93.8 E10, and with NuMI, and MCenter taking beam.

Operations put the Recycler into shot setup at 2:13 AM to stash antiprotons.

Pbar returns to stacking at 3:19 AM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 7:41 AM to stash antiprotons.

Pbar returns to stacking at 9:00 AM, MCenter is off for analyzing magnet repair.

Operations put the Recycler into shot setup at 12:44 PM to stash antiprotons.

Pbar returns to stacking at 2:21 PM.

At 5:18 PM, MiniBooNE began taking beam once again.

MCenter turned back on at 7:04 PM, they are running without their analyzing magnets. Work will continue on MC7AN1 tomorrow.

Operations put the Recycler into shot setup at 7:26 PM to stash antiprotons.

Pbar returns to stacking at 8:44 PM, beam also returns to MCenter.

Friday January 13

The midnight shift began with Operations monitoring store #4589, the Recycler Ring (RR) stash at 254.7 E10, Pbar stack at 67.6 E10, and with NuMI, MiniBooNE, and MCenter taking beam.

Operations put the Recycler into shot setup at 1:04 AM to stash antiprotons.

Pbar returns to stacking at 1:51 AM, beam also returns to MCenter.

Store #4589 was terminated at 4:03 AM.

Shot setup begins at 4:22 AM.

MS6LCW water system tripped off at 6:36 AM, experts are investigating.

Operations established store 4590 at 7:36 AM with an initial luminosity of 161.36 E30.

MCenter turned off for analyzing magnet repair at 7:37 AM

Pbar returns to stacking at 7:47 AM.

The Plan for Friday

Keep the store, access Pbar Rings/Transport and Booster for necessary repairs. Then stack, stash and store over the long weekend.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/13/2006   email Fermilab